# MAGID: An Automated Pipeline for Generating Synthetic Multi-modal Datasets

### v1

@@ Abstract
Development of multimodal interactive systems is hindered by the lack of rich, multimodal (text, images) conversational data, which is needed in large quantities for LLMs. Previous approaches augment textual dialogues with retrieved images, posing privacy, diversity, and quality constraints. In this work, we introduce Multimodal Augmented Generative Images Dialogues (MAGID), a framework to augment text-only dialogues with diverse and high-quality images  [^1] The link to code: http://anon_for_review.com. Subsequently, a diffusion model is applied to craft corresponding images, ensuring alignment with the identified text. Finally, MAGID incorporates an innovative feedback loop between an image description generation module (textual LLM) and image quality modules (addressing aesthetics, image-text matching, and safety), that work in tandem to generate high-quality and multi-modal dialogues. We compare MAGID to other SOTA baselines on three dialogue datasets, using automated and human evaluation. Our results show that MAGID is comparable to or better than baselines, with significant improvements in human evaluation, especially against retrieval baselines where the image database is small.
@@ 3.4.1 Feedback Loop
Model
Accuracy
Precision
Recall
F1 score
CLIP score
MM-Relevance
Aesthetic
#images
GPT 4
67.24%
70.49%
46.87%
0.56 0.27 294.52
0.57
1359
GPT 3.5
63.54%
69.43%
33.97%
0.46
0.26
293.51
0.58
1001
Falcon-40b-Ins.
58.93%
61.26%
24.13%
0.35
0.25
254.50
0.58
794
Koala 13b
56.28%
62.33%
6.91%
0.12
0.25
243.31
0.57
223
Llama 13b
57.10%
60.00%
13.64%
0.22
0.25
247.99
0.57
460
OpenLLaMA
57.94%
64.36%
12.69%
0.21
0.25
250.96
0.58
390
Vicuna 13b
58.77%
66.60%
14.38%
0.24
0.26
255.18
0.57
506
MMDialogue [^3] Ground Truth
N/A
N/A
N/A
N/A
0.262
N/A
0.47
2717
@@ 4.1 Quantitative Evaluation
Mean MAGID
Mean MMDD
Gwet’s AC1
Mean MAGID
Mean MMDial.
Gwet’s AC1
Mean MAGID
Mean Photo.
Gwet’s AC1
Q1
96.29%
3.71%
0.74
48.17%
51.83%
0.63
58.11%
41.89%
0.47
Q2
96.29%
3.71%
0.89
49.33%
50.67%
0.65
68.24%
31.76%
0.71
Q3
89.11%
10.89%
0.75
52.72%
47.28%
0.54
64.90%
35.10%
0.53
Q4
91.11%
8.89%
0.83
46.31%
53.69%
0.65
61.98%
38.02%
0.54
Q5
95.57%
4.43%
0.89
51.94%
48.06%
0.63
64.02%
35.98%
0.61
Q6
80.92%
19.08%
0.65
63.90%
36.10%
0.55
69.99%
30.01%
0.64
@@
Prompt
Accuracy
Precision
Recall
F1 score
ZS
65.53%
73.12%
36.16%
0.48
FS
63.89%
69.67%
34.45%
0.46
CoT
68.51%
73.37%
47.32%
0.57
@@ 6. Q6:
Model
Accuracy
Precision
Recall
F1 score
CLIP score
MM-Relevance
Aesthetic
#images
GPT 4
67.24%
70.49%
46.87%
0.56 0.27 294.52
0.57
1359
GPT 3.5
63.54%
69.43%
33.97%
0.46
0.26
293.51
0.58
1001
Falcon-40b-Ins.
58.93%
61.26%
24.13%
0.35
0.25
254.50
0.58
794
OpenLLaMA
57.94%
64.36%
12.69%
0.21
0.25
250.96
0.58
390
GPT 4
67.86%
69.70%
50.64%
0.59 0.27 282.25
0.55
1485
GPT 3.5
68.51%
73.37%
47.32%
0.57
0.26
278.16
0.55
1109
Falcon-40b-Ins.
56.77%
53.58%
28.80%
0.37
0.23
224.59
0.55
1075
OpenLLaMA
58.92%
62.50%
21.51%
0.32
0.21
213.56
0.56
696
@@ 5 MAGID Dataset
As a proof of concept, and consistent with studies like Lee et al. (2021), we employed text-only datasets such as DailyDialog Li et al. (2017), Persona-Chat Zhang et al. (2018), and PhotoChat Zang et al. (2021) (by replacing its images with MAGID) to generate a multi-modal dataset  [^4] The link to dataset: http://anon_for_review.com of 53,620 dialogues. Based on the results of our experiments, we used GPT-3.5 to transform 47,868 input dialogues and GPT-4 to augment the rest. Table 5 shows the statistics of the generated dataset with MAGID. The data and the code will be made available to the public upon acceptance.
@@
 Category
Train
Test
Total dialogues
47643
5977
Avg length of dialogues
11.76
11.36
Avg length of sentences
9.77
9.60
Total images
67951
10229
@@ Appendix B PhotoChat results
Model
Accuracy
Precision
Recall
F1 score
CLIP score
MM-Relevance
#images
Aesthetic
GPT 3.5
86.11%
28.62%
25.91%
0.27
0.25
313.64
87
0.57
Falcon-40b-Ins.
88.10%
28.04%
11.83%
0.17
0.24
303.68
403
0.58
Koala 13b
89.61%
30.43%
2.94%
0.05
0.24
283.44
92
0.61
Llama 13b
87.32%
20.79%
9.54%
0.13
0.23
244.36
433
0.59
OpenLLaMA
88.75%
27.31%
8.03%
0.12
0.23
270.36
696
0.59
Vicuna 13b
88.40%
25.48%
8.35%
0.13
0.24
244.97
602
0.55
PhotoChat
N/A
N/A
N/A
N/A
0.213
N/A
961
0.49
@@ Appendix C Image Generator Ablation Study
 Model
Aesthetic Score
CLIP Score
SDXL 1.0
0.56
0.26
SDXL 0.9
0.57
0.26
Stable Diffusion 2.0
0.53
0.26
@@ Appendix E Downstream Training
 Dataset
PPL
BLEU-1
BLEU-2
distinct-1
distinct-2
MMDialogue
73.09
8.3 3.9
0.94
0.965
MAGID
70.99
7.9
3.9
0.94
0.971
MMDD
78.86
7.5
3.0
0.93
0.963
No image
78.88
7.9
3.0
0.92
0.952

### v2

@@ Abstract
Development of multimodal interactive systems is hindered by the lack of rich, multimodal (text, images) conversational data, which is needed in large quantities for LLMs. Previous approaches augment textual dialogues with retrieved images, posing privacy, diversity, and quality constraints. In this work, we introduce Multimodal Augmented Generative Images Dialogues (MAGID), a framework to augment text-only dialogues with diverse and high-quality images 111code link: https://github.com/amazon-science/MAGID. Subsequently, a diffusion model is applied to craft corresponding images, ensuring alignment with the identified text. Finally, MAGID incorporates an innovative feedback loop between an image description generation module (textual LLM) and image quality modules (addressing aesthetics, image-text matching, and safety), that work in tandem to generate high-quality and multi-modal dialogues. We compare MAGID to other SOTA baselines on three dialogue datasets, using automated and human evaluation. Our results show that MAGID is comparable to or better than baselines, with significant improvements in human evaluation, especially against retrieval baselines where the image database is small.
@@ 3.4.1 Feedback Loop
Model Accuracy Precision Recall F1 score CLIP score MM-Relevance Aesthetic #images
GPT 4 67.24% 70.49% 46.87% 0.56 0.27 294.52 0.57 1359
GPT 3.5 63.54% 69.43% 33.97% 0.46 0.26 293.51 0.58 1001
Falcon-40b-Ins. 58.93% 61.26% 24.13% 0.35 0.25 254.50 0.58 794
Koala 13b 56.28% 62.33% 6.91% 0.12 0.25 243.31 0.57 223
Llama 13b 57.10% 60.00% 13.64% 0.22 0.25 247.99 0.57 460
OpenLLaMA 57.94% 64.36% 12.69% 0.21 0.25 250.96 0.58 390
Vicuna 13b 58.77% 66.60% 14.38% 0.24 0.26 255.18 0.57 506
MMDialogue [^3] Ground Truth N/A N/A N/A N/A 0.262 N/A 0.47 2717
@@ 4.1 Quantitative Evaluation
# Mean MAGID Mean MMDD Gwet’s AC1 Mean MAGID Mean MMDial. Gwet’s AC1 Mean MAGID Mean Photo. Gwet’s AC1
Q1 96.29% 3.71% 0.74 48.17% 51.83% 0.63 58.11% 41.89% 0.47
Q2 96.29% 3.71% 0.89 49.33% 50.67% 0.65 68.24% 31.76% 0.71
Q3 89.11% 10.89% 0.75 52.72% 47.28% 0.54 64.90% 35.10% 0.53
Q4 91.11% 8.89% 0.83 46.31% 53.69% 0.65 61.98% 38.02% 0.54
Q5 95.57% 4.43% 0.89 51.94% 48.06% 0.63 64.02% 35.98% 0.61
Q6 80.92% 19.08% 0.65 63.90% 36.10% 0.55 69.99% 30.01% 0.64
@@
Prompt Accuracy Precision Recall F1 score
ZS  65.53% 73.12% 36.16% 0.48
FS  63.89% 69.67% 34.45% 0.46
CoT  68.51% 73.37% 47.32% 0.57
@@ 6. Q6:
Model Accuracy Precision Recall F1 score CLIP score MM-Relevance Aesthetic #images
GPT 4 67.24% 70.49% 46.87% 0.56 0.27 294.52 0.57 1359
GPT 3.5 63.54% 69.43% 33.97% 0.46 0.26 293.51 0.58 1001
Falcon-40b-Ins. 58.93% 61.26% 24.13% 0.35 0.25 254.50 0.58 794
OpenLLaMA 57.94% 64.36% 12.69% 0.21 0.25 250.96 0.58 390
GPT 4 67.86% 69.70% 50.64% 0.59 0.27 282.25 0.55 1485
GPT 3.5 68.51% 73.37% 47.32% 0.57 0.26 278.16 0.55 1109
Falcon-40b-Ins. 56.77% 53.58% 28.80% 0.37 0.23 224.59 0.55 1075
OpenLLaMA 58.92% 62.50% 21.51% 0.32 0.21 213.56 0.56 696
@@ 5 MAGID Dataset
As a proof of concept, and consistent with studies like Lee et al. (2021), we employed text-only datasets such as DailyDialog Li et al. (2017), Persona-Chat Zhang et al. (2018), and PhotoChat Zang et al. (2021) (by replacing its images with MAGID) to generate a multi-modal dataset  [^4] The link to dataset: https://github.com/amazon-science/MAGID of 53,620 dialogues. Based on the results of our experiments, we used GPT-3.5 to transform 47,868 input dialogues and GPT-4 to augment the rest. Table 5 shows the statistics of the generated dataset with MAGID. The data and the code will be made available to the public upon acceptance.
@@
 Category Train Test
Total dialogues  47643 5977
Avg length of dialogues  11.76 11.36
Avg length of sentences  9.77 9.60
Total images  67951 10229
@@ Appendix B PhotoChat results
Model Accuracy Precision Recall F1 score CLIP score MM-Relevance #images Aesthetic
GPT 3.5 86.11% 28.62% 25.91% 0.27 0.25 313.64 87 0.57
Falcon-40b-Ins. 88.10% 28.04% 11.83% 0.17 0.24 303.68 403 0.58
Koala 13b 89.61% 30.43% 2.94% 0.05 0.24 283.44 92 0.61
Llama 13b 87.32% 20.79% 9.54% 0.13 0.23 244.36 433 0.59
OpenLLaMA 88.75% 27.31% 8.03% 0.12 0.23 270.36 696 0.59
Vicuna 13b 88.40% 25.48% 8.35% 0.13 0.24 244.97 602 0.55
PhotoChat N/A N/A N/A N/A 0.213 N/A 961 0.49
@@ Appendix C Image Generator Ablation Study
 Model Aesthetic Score CLIP Score
SDXL 1.0  0.56 0.26
SDXL 0.9  0.57 0.26
Stable Diffusion 2.0  0.53 0.26
@@ Appendix E Downstream Training
 Dataset PPL BLEU-1 BLEU-2 distinct-1 distinct-2
MMDialogue  73.09 8.3 3.9 0.94 0.965
MAGID  70.99 7.9 3.9 0.94 0.971
MMDD  78.86 7.5 3.0 0.93 0.963
No image  78.88 7.9 3.0 0.92 0.952